OpenStudy (anonymous):

How many kilograms of pure salt must be added to 7 kg of a 30% salt solution in order to produce a 60% solution?

OpenStudy (rulnick):

( 2.1kg + x kg ) / (7 kg + x kg) = 60/100 210 + 100x = 420 + 60x 40x = 210 x = 21/4 or 5.25 answer: 5.25 kg check: 5.25+2.1=7.35, 7.35/(7+5.25) = 7.35/12.25 = 0.6 = 60%

OpenStudy (rulnick):

In case you are wondering, the 2.1 kg is the amount of salt in the original 7 kg solution (30% of 7 kg = 2.1 kg).